# Model Alignment Search

### v5

@@ Abstract
When can we say that two neural systems are the same? The answer to this question is goal-dependent, and it is often addressed through correlative methods such as Representational Similarity Analysis (RSA) and Centered Kernel Alignment (CKA). What nuances do we miss, however, when we fail to causally probe the representations? Do the dangers of cause vs. correlation exist in comparative representational analyses? In this work, we introduce a method for connecting neural representational similarity to behavior through causal interventions. The method learns orthogonal transformations that find an aligned subspace in which behavioral information from multiple distributed networks’ representations can be isolated and interchanged. We first show that the method can be used to transfer the behavior from one frozen Neural Network (NN) to another in a manner similar to model stitching, and we show how the method can complement correlative similarity measures like RSA. We then introduce an efficient subspace orthogonalization technique using the Gram-Schmidt process—that can also be used for Distributed Alignment Search (DAS)—allowing us to perform analyses on larger models. Next, we empirically and theoretically show how our method can be equivalent to model stitching when desired, or it can take a form that is more restrictive to causal information, and in both cases, it reduces the number of required matrices for a comparison of n models from quadratic to linear in n. We then show how we can augment the loss objective with an auxiliary loss to train causally relevant alignments even when we can only read the representations from one of the two networks during training (like with biological networks). Lastly, we use number representations as a case study to explore how our method can be used to compare specific types of representational information across tasks and models.
@@ 1 Introduction
An important question for understanding both Artificial and Biological Neural Networks (ANNs and BNNs) is knowing what it means for one distributed system to model or represent another (sucholutsky2023repalign, ). Establishing isomorphisms between different distributed systems can be useful for simplifying their complexity and for understanding otherwise opaque inner mechanisms. We cannot yet measure from every individual neuron in most BNNs; even if we could, as is the case in ANNs, it is still difficult to find satisfying ways of understanding the neural behavior. Finding simplified models that exhibit the causal relationships of more complex distributed systems can make complex systems more interpretable and communicable, potentially leading to useful insights (cao_explanatory_2021, ; cao_explanatory_2024, ; richards2019dlframework, ). Furthermore, there are a number of open questions about how representations differ or converge between architectures, tasks, and modalities (huh2024platonic, ; sucholutsky2023repalign, ; wang2024universality, ; li_vision_2024, ; Hosseini2024universality, ; zhang2024assessinglearningalignmentunimodal, ; grant2024das, ).
@@
Figure 1: (a) A depiction of an interchange intervention on the target latent vector $h_{t,trg}^{(i)}$ at time $t$ from model $\psi_{i}$ using the source latent vector $h_{u,src}^{(j)}$ at time $u$ from $\psi_{j}$. Rectangles represent vectors; colors depict different types of encoded information. The causally relevant information is spread across the neural population, represented by the red and green semi-vertical slices in the respective $h$ vectors. To disentangle and isolate the causal information, we rotate the $h$ vectors into an aligned space, using $Q_{i}$ and $Q_{j}$, where the relevant information is organized orthogonally to the extraneous information. We can then transfer the relevant information without affecting other types of information by applying binary masks (black represents 0s and white 1s and $\odot$ is a Hadamard product) to the $z$ vectors and taking their sum. We then invert the rotation on $z_{t,v}^{(i)}$ to return it to the original latent space where it can be used by $\psi_{i}$. (b) Depicts Stepwise MAS where the individual intervention shown in panel (a) is applied at multiple token positions in the sequence. We limit our interventions to contiguous sets of tokens starting from the first token.
@@
Researchers often use correlational methods to measure the similarity of different neural representations. We can see examples of this in works that perform direct correlational analyses between individual ANN activations and BNN firing rates (Yamins2016, ; Maheswaranathan2019, ; khosla2023softmatching, ; williams2022procrustes, ), and in works that use Representational Similarity Analysis (RSA)—or Centered Kernel Alignment (CKA) (kornblith2019cka, ; williams2024rsacka, )—finding 2nd order isomorphisms between model and system (kriegeskorte2008rsa, ). We also see examples of this in linear decoding techniques, where linear decodability and predictability can be used as a measure of the type of information encoded in distributed representations (chen2020simclr, ; radford2021clip, ; grill2020byol, ; caron2021dino, ; Haxby2001ogmvpa, ; Haxby2013mvpareview, ; lampinen2024learned, ). A question remains, however, how to causally associate these neural analyses with behavioral outcomes. What do we miss when we ignore the behavioral relevance of the neural activity (geiger2024causalabstractiontheoreticalfoundation, ; cloos2024differentiableoptimizationsimilarityscores, ; schaeffer2024doesregressbrain, )?
@@
Some works have made progress toward understanding behavioral representation similarity by transforming intermediate representations from one system into a usable form for a model. We see examples in works like sexton_reassessing_2022 where they attempt to use transformed neural recordings in a trained computational model, and in model stitching, where a linear mapping is learned from intermediate representations from one ANN to another for the purpose of measuring similarity or improving one of the two models (lahner2023directalignment, ; moschella_relative_2023, ; bansal2021modelstitching, ; lenc2015modelstitching, ). To build upon these works, we ask: 1) what do these representational mappings tell us about the underlying representations of the two systems; are behaviorally successful mappings both necessary and sufficient for claims of functional similarity? 2) How can we improve the efficiency of existing methods, and in the process, how can we isolate the behaviorally relevant activity within the representations? 3) How do we achieve causal relevance in systems that we do not have causal access—as is often the case in ANN to BNN comparisons? And 4) How do we isolate the similarity of specific behavioral information in the representations, thus allowing us to compare models with disparate behavioral outputs?
@@
In this work, we introduce a method, Model Alignment Search (MAS), that uses causal interventions to isolate and compare behaviorally relevant activity from neural representations from different ANNs. MAS can be thought of as an efficient form of model stitching (lenc2015modelstitching, ; bansal2021modelstitching, ) or a multi-model extension of Distributed Alignment Search (DAS) (geiger2021, ; geiger2023boundless, ; grant2024das, )–––a technique used to align ANNs to directed acyclic graphs. MAS learns a rotation matrix for each model in the comparison with the goal of finding an aligned subspace where behavioral information can be causally interchanged between and within the models. These manipulated representations are then returned back to their original neural space where they are used to produce behavior that can be compared to expected behavior.
@@
    We introduce and motivate MAS by comparing to correlative representational similarity methods.
@@
    We introduce an efficient orthogonal intervention method based on the Gram-Schmidt process that can be used to isolate functional subspaces in both MAS and DAS on large representations.
@@
    We empirically and theoretically compare MAS to model stitching, showing that it is more efficient when comparing 3 or more models, and it can be more restrictive than model stitching, making it a better choice for some representational analyses.
@@
    We introduce a counterfactual latent auxiliary loss objective that can be used to find behaviorally relevant alignments in cases where one of the two models is causally inaccessible (making the future of the technique relevant for comparisons between ANNs and BNNs).
@@ 2 Methods
2.1 Tasks
Each task consists of a sequence of tokens that start with a beginning of sequence token, B, and end with an end of sequence token, E. Some of the tokens in each task are produced by the task environment and define the specific goals within the task. Other tokens are determined by these task provided tokens. Each trial is considered correct when all deterministic tokens are correctly predicted. During the model training, we include all token types in a NTP cross entropy loss, even though many tokens are unpredictable.
@@ 2.1 Tasks
To prevent solutions that use positional readout (grant2024das, ), we modify the task for transformer trainings: each token in the demo phase has a 0.2 probability of being a "void" token type, V, that has no impact on the object quantity. An example sequence with an object count of 2 could be: "B V $\text{D}_{b}$ V V $\text{D}_{c}$ T R R E". All evaluations and analyses use the original Multi-Object Task.
@@ 2.2 Model Architectures
Each model in this work is autoregressively trained to perform only one of the tasks through NTP. We train 2 model seeds for each task variant up to $>99.99\%$ accuracy on both training and validation data and freeze the weights before analysis and interpretation. We consider Gated Recurrent Units (GRUs) (Cho2014GRU, ), Long-Short Term Memory recurrent networks (LSTMs) (hochreiter1997lstm, ), and two layer Transformers based on the Roformer architecture (vaswani2017, ; touvron2023llama, ; su2023roformer, ). We also include one analysis using DeepSeek R1 Distill Llama 8B guo2025deepseek . The custom GRUs and Transformers use a dimensionality of 128, whereas the LSTM uses 64 dimensions for each the h and c vectors. We leave details of GRU and LSTM cells to the referenced papers beyond noting that the GRU and LSTM based models have the structure:
$h_{t+1}$ $=f(h_{t},x_{t})$ (1)
$\hat{x}_{t+1}$ $=g(h_{t+1})$ (2)
Where $h_{t}$ is the hidden state vector at step $t$, $x_{t}$ is the input token at step $t$, $f$ is the recurrent function (either a GRU or LSTM cell), and $g$ is a two layer (two matrix) feed-forward network (FFN) used to make a prediction, $\hat{x}_{t+1}$, of the token at step $t+1$ from the updated hidden state $h_{t+1}$.
The transformer architecture uses Rotary Positional Encodings (RoPE) (su2023roformer, ) and GELU nonlinearities (hendrycks2023gelus, ). Transformers use a history of input tokens, $X_{t}=[x_{1},x_{2},...,x_{t}]$, at each time step, $t$, to make a prediction: $\hat{x}_{t+1}=f(X_{t})$, where $f$ is now the transformer architecture. We show results from 2 layer, single attention head transformers. We refer readers to Figure 4 for more details. See training details in Appendix A.1. We sometimes refer to the NNs generally as $\psi$.

### v6

@@ Abstract
When can we say that two neural systems are the same? The answer to this question is goal-dependent, and it is often addressed through correlative methods such as Representational Similarity Analysis (RSA) and Centered Kernel Alignment (CKA). What nuances do we miss, however, when we fail to causally probe the representations of the systems? Do the dangers of cause vs. correlation exist in comparative representational analyses? In this work, we introduce a method for connecting neural representational similarity to behavior through causal interventions. The method learns orthogonal transformations that find an aligned subspace in which behavioral information can be isolated and interchanged between multiple distributed networks’ representations. We first show that the method can be used to transfer the behavior from one frozen Neural Network (NN) to another in a manner similar to model stitching, and we show how the method can differ from correlative similarity measures like RSA. Next, we empirically and theoretically show how our method can be equivalent to model stitching when desired, or it can take a form that is more restrictive, narrowing its focus to shared causal information; and in both forms, it reduces the number of required matrices for a comparison of n models to be linear in n. We then present a case study in number-related tasks showing that the method can be used to examine specific subtypes of causal information, and we present another case study showing that the method can reveal toxicity in fine-tuned DeepSeek-r1-Qwen-1.5B models. Lastly, we show how we can augment the loss with a counterfactual latent auxiliary objective to improve causal relevance when one of the two networks during training is causally inaccessible (as is often the case in comparisons with biological networks). We use our results to encourage the use of causal methods in neural similarity analyses and to suggest future explorations of network similarity methodology for understanding LLM misalignment.
@@ 1 Introduction
An important question for understanding both Artificial and Biological Neural Networks (ANNs and BNNs) is knowing what it means for one distributed system to model or represent another (Sucholutsky et al., 2023). Establishing isomorphisms between different distributed systems can be useful for simplifying their complexity and for understanding otherwise opaque inner mechanisms. Instead of asking how to interpret neural activity directly, we can instead use simplified neural systems, or neural systems understood through a higher level of analysis, and compare the internals of these known systems to uknown systems (Cao & Yamins, 2021; 2024; Richards et al., 2019). Furthermore, there are a number of open questions about how representations differ or converge between architectures, tasks, and modalities (Huh et al., 2024; Sucholutsky et al., 2023; Wang et al., 2024; Li et al., 2024; Hosseini et al., 2024; Zhang et al., 2024; Grant et al., 2025). Implicitly intertwined in these issues is the notion of network similarity, or dissimilarity. How do we measure the degree to which one model represents another?
@@
Figure 1: (a) A depiction of an interchange intervention from Equation 6 on the target latent vector $h^{(trg)}_{\psi_{i}}$ from model $\psi_{i}$ using the source latent vector $h^{(src)}_{\psi_{k}}$ from $\psi_{k}$. Rectangles represent vectors; colors distinguish between behaviorally relevant and extraneous activity. The causally relevant information is spread across the neural population (Smolensky, 1988; Park et al., 2023), represented by the red and green semi-vertical slices in the respective $h$ vectors. To disentangle and isolate the complement of the behavioral null-space, we rotate the $h$ vectors into an aligned space, using $Q_{\psi_{i}}$ and $Q_{\psi_{k}}$, where the behaviorally relevant information is organized along separate dimensions than the extraneous, behavioral null-space. We can then intervene on and transfer the relevant information without affecting other information. In the figure, this is done by applying binary masks (black represents 0s and white 1s and $\odot$ is a Hadamard product) to the $z$ vectors and taking their sum. We then invert the rotation on $z^{(v)}_{\psi_{i}}$ to return it to the original latent space where it can be used by $\psi_{i}$ to make predictions. (b) Depicts Stepwise MAS where the individual intervention shown in panel (a) is applied at multiple token positions in the sequence. We limit our interventions to contiguous sets of tokens starting from the first token and ending with a sampled position $t$.
@@
Researchers often use correlational methods to measure the similarity of different neural representations. We can see examples of this in works that perform direct correlational analyses between individual ANN activations and BNN firing rates (Yamins & DiCarlo, 2016; Maheswaranathan et al., 2019; Khosla & Williams, 2023; Williams et al., 2022), and in works that use Representational Similarity Analysis (RSA)—or Centered Kernel Alignment (CKA) (Kornblith et al., 2019; Williams, 2024)—finding 2nd order isomorphisms between model and system (Kriegeskorte et al., 2008). We also see examples of this in linear decoding techniques, where linear decodability and predictability can be used as a measure of the type of information encoded in distributed representations (Chen et al., 2020; Radford et al., 2021; Grill et al., 2020; Caron et al., 2021; Haxby et al., 2001; Haxby, 2013; Lampinen et al., 2024). A question remains, however, how to causally associate these neural analyses with behavioral outcomes. What do we miss when we ignore the behavioral relevance of the neural activity (Pearl, 2010; Geiger et al., 2024; Cloos et al., 2024; Schaeffer et al., 2024)? And how do we incorporate a focus on behavioral relevance in our measures of neural similarity?
@@
Some works have made progress towards connecting representational similarity to behavior by transforming intermediate representations from one neural system into a usable form for another. We see examples in works like Sexton & Love (2022) where they attempt to use transformed neural recordings in a trained computational model, and in model stitching, where a linear mapping is learned from intermediate representations from one ANN to another for the purpose of measuring similarity and/or improving one of the two models (Lähner & Moeller, 2024; Moschella et al., 2023; Bansal et al., 2021; Lenc & Vedaldi, 2015). To build upon these works, we ask: 1) what do these representational mappings tell us about the underlying representations of the two systems; are behaviorally successful mappings both necessary and sufficient for claims of functional similarity? 2) How can we isolate and compare the behaviorally relevant activity within transferred representations? 3) How do we isolate the similarity of specific behavioral information in the representations, thus allowing us to compare models with disparate behavioral outputs? And 4) how do we achieve causal relevance in systems for which we do not have causal access—as is often the case in ANN to BNN comparisons?
@@
In this work, we introduce a neural similarity method, Model Alignment Search (MAS), that uses causal interventions to isolate and compare behaviorally relevant activity from neural representations in different ANNs. MAS learns a rotation matrix, or more generally an alignment function, for each model in the analysis to simultaneously uncover causally relevant latent subspaces in each model and map these spaces onto one another so that behavioral information can be patched between models. These manipulated representations are then returned back to their original neural space where the networks’ resulting behavior can be compared to the expected counterfactual behavior as a means of measuring neuro-behavioral similarity.
@@
    We introduce and motivate MAS by comparing to correlative representational similarity methods demonstrating a need to supplement correlational with causal methods depending on the research goals.
@@
    We empirically and theoretically compare MAS to model stitching, showing that it requires less compute when comparing 3 or more models, and it can be more restrictive to causal subspaces than model stitching.
@@
    Using model misalignment as a case study, we show that the method can be used to reveal toxicity in DeepSeek-R1-Qwen-1.5B models (Guo et al., 2025) finetuned to be toxic or nontoxic.
@@
    We introduce a counterfactual latent auxiliary loss objective that can be used to find behaviorally relevant alignments in cases where one of the two models is causally inaccessible (making the technique potentially relevant for comparisons between ANNs and BNNs).
2 Background and Related Work
There has been a call to use causal methods for both interpreting neural activity and making comparisons between different NN internals (Geiger et al., 2024; Sexton & Love, 2022; Feather et al., 2025; Lampinen et al., 2025). This call has been made due to principles of causal mediation Pearl (2010), and, more recently, representation and causal computation have been shown to be different in NNs (Hermann & Lampinen, 2020; Braun et al., 2025; Lampinen et al., 2025). In the mechanistic interpretability literature, a popular method for making claims about cause and effect in NNs is activation patching (Geiger et al., 2020; Vig et al., 2020; Wang et al., 2022; Meng et al., 2023) where experimentalists manipulate the NN activity to observe downstream effects. One such method is DAS, which finds causal subspaces that align with high-level variables from causal abstractions (Geiger et al., 2021; 2023; Wu et al., 2024). In cases of multi-model comparisons, a particularly popular method for causally comparing representations between different models is model stitching (Lenc & Vedaldi, 2015; Bansal et al., 2021). In this section, we present a background on both model stitching and DAS to establish notation and a conceptual framework for MAS.
2.1 Model Stitching
Model stitching aims to learn a mapping from the latent state space of one NN, $\psi_{1}$, to that of another, $\psi_{2}$. Let $h_{\psi_{i}}\in\mathbb{R}^{d_{\psi_{i}}}$ denote a latent representation from $\psi_{i};\,i\in\{1,2\}$. Stitching seeks a transformation $T:\mathbb{R}^{d_{\psi_{1}}}\to\mathbb{R}^{d_{\psi_{2}}}$ such that
$T\!\left(h_{\psi_{1}}\right)\approx h_{\psi_{2}}$
where $h_{\psi_{2}}\in\mathbb{R}^{d_{\psi_{2}}}$ is the corresponding representation in $\psi_{2}$. The mapped representation $T(h_{\psi_{1}})$ is then fed into $\psi_{2}$’s remaining layers, denoted $f_{\psi_{2}}(x,h)$, so that $\psi_{2}$ can make predictions consistent with those produced by $\psi_{1}$.
Although $T$ can be trained in a number of ways, we focus on cases in which $T$ is learned by optimizing $\psi_{2}$’s outputs, conditioned on the stitched representation $\hat{h}_{\psi_{2}}=T(h_{\psi_{1}})$, to match $\psi_{1}$’s behavior, also conditioned on $h_{\psi_{1}}$. The learning objective is a behavioral matching loss:
$\mathcal{L}(T)\;=\;\mathbb{E}_{x\sim\mathcal{D}}\big{[}\ell\big{(}f_{\psi_{1}}(x,h_{\psi_{1}})\,\|\,f_{\psi_{2}}(x,\hat{h}_{\psi_{2}})\big{)}\big{]}$ (1)
where $x~\mathcal{D}$ are samples from a dataset and $\ell$ is a divergence measure (e.g., KL divergence or cross-entropy) between $\psi_{1}$’s output distribution and $\psi_{2}$’s output distribution under the stitched representation.
In this work, we narrow our focus to transformations of the form
$T(h_{\psi_{1}})=aWh_{\psi_{1}}$
where $a\in\mathbb{R}$ is a scalar constant and $W\in\mathbb{R}^{d_{\psi_{2}}\times d_{\psi_{1}}}$ is a matrix—possibly low rank—with orthonormal rows.
Model stitching has been used as a method to make claims about the similarity between two different NNs Bansal et al. (2021). However, it has problems as a similarity method due to the fact that it is unidirectional, and, as we will show in Section 4.2, it can make use of $\psi_{1}^{\prime}s$ behavioral null space when mapping to $\psi_{2}$—where the behavioral null space is defined as the span of vector directions that have no influence on the NN’s behavior. Furthermore, a successful mapping ignores naturally occurring variability in $\psi_{2}$’s behavioral null space. Thus, a successful mapping $T$ only needs to produce a single sufficient representation for any given behavior. Thus, successful $T$’s do not necessarily indicate that $\psi_{1}$ and $\psi_{2}$ perform the task in the same way.
2.2 DAS Formulation
DAS is a framework for causally determining the degree of alignment between a single NN’s latent representations and variables from a causal abstraction (Geiger et al., 2021; 2023; Wu et al., 2024). We can think of such variables as those in a computer program, or nodes in a Directed Acyclic Graph (DAG). DAS does this alignment by attempting to transform model latent states $h\in\mathbb{R}^{d_{\psi}}$ into aligned vectors $z\in\mathbb{R}^{d_{\psi}}$ that consist of contiguous subspaces that have an analogous causal effect on behavior as high-level variables from a pre-specified causal abstraction. The transformation is performed by learnable, invertible Alignment Function (AF), $\mathcal{A}$, as follows: $z=\mathcal{A}(h)$ (Grant et al., 2025). We narrow our focus to orthogonal AFs: $\mathcal{A}(h)=Qh$ where $Q\in\mathbb{R}^{d_{\psi}\times d_{\psi}}$ is an orthogonal matrix. The purpose of this transformation is that it organizes causal subspaces into separated, contiguous dimensions in $z$. This allows for causal interventions on the values of specific subspaces—analogously to changing value of a variable in a computer program—without affecting the values of the other subspaces.
Concretely, for a given causal abstraction with variables $\text{var}_{i}\in\{\text{var}_{1},\text{var}_{2},...,\text{var}_{n}\}$, DAS learns a rotation $Q$ to find a $z$ composed of subspaces $\vec{z}_{\text{var}_{i}}\in\mathbb{R}^{d_{\text{var}_{i}}}$ corresponding to each of the variables of the causal abstraction. We also include a causally irrelevant subspace, $\vec{z}_{\text{extra}}\in\mathbb{R}^{d_{\text{extra}}}$, encoding extraneous, behaviorally irrelevant activity (i.e., the behavioral null-space and dormant subspaces—defined as those that do not vary between inputs (Makelov et al., 2023)).
$Q(h)=z=\begin{bmatrix}\vec{z}_{\text{var}_{1}}\\ \vec{z}_{\text{var}_{2}}\\ \cdots\\ \vec{z}_{\text{var}_{n}}\\ \vec{z}_{\text{extra}}\end{bmatrix}$ (2)
Each $\vec{z}_{\text{var}_{i}}\in\mathbb{R}^{d_{\text{var}_{i}}}$ is a column vector of potentially different lengths satisfying the relation $d_{\text{extra}}+\sum_{i=1}^{n}d_{\text{var}_{i}}=d_{\psi}$. Under this assumption, the value of a single causal variable encoded in $h$ can be freely exchanged by performing an interchange intervention defined as follows:
$h^{(v)}=Q^{-1}((1-D_{\text{var}_{i}})Qh^{(trg)}+D_{\text{var}_{i}}Qh^{(src)})$ (3)
Where $D_{\text{var}_{i}}\in\mathbb{R}^{d_{\psi}\times d_{\psi}}$ is a manually chosen, diagonal, binary matrix with $d_{\text{var}_{i}}$ contiguous ones along the diagonal used to isolate the dimensions that make up $\vec{z}_{\text{var}_{i}}$, $h^{(src)}$ is the source vector from which the subspace activity is harvested, $h^{(trg)}$ is the target vector into which the harvested activations are substituted, and $h^{(v)}$ is the resulting intervened vector that is used to replace $h^{(trg)}$ in the model’s processing. This allows the model to make predictions using a different value of subspace $\text{var}_{i}$ after a successful intervention.
DAS uses counterfactual behavior to create intervention data to train and evaluate $Q$. Counterfactual behavior is defined as the behavior that would have occurred had the value of a variable in a causal abstraction been different and everything else remained the same (Pearl, 2010; Geiger et al., 2021). We can create intervention data by freezing the state of an environment at a particular point in time, changing one or more values in the causal abstraction’s variables, and then unfreezing the environment and using the causal abstraction to generate the counterfactual behavior. These intervention samples can then be used as training labels to train $Q$ while keeping the NN parameters frozen. Once $Q$’s training has converged, the robustness of the alignment can be evaluated using the Interchange Intervention Accuracy (IIA)—defined as the NN’s prediction accuracy on the counterfactual behavior for a held out set of interventions. IIA is then used to make claims about the NN’s internal mechanisms.
@@ 3 Methods
All of the analyses in this work are performed on autoregressive, sequence based ANNs trained to predict sequences of tokens. Depending on the experiment, we consider GRUs (Cho et al., 2014), LSTMs (Hochreiter & Schmidhuber, 1997), shallow transformers (Vaswani et al., 2017; Su et al., 2023), and DeepSeek-R1-Distill-Qwen-1.5B (Guo et al., 2025). The majority of our analyses are confined to models trained to perform variations of numeric tasks. These tasks serve as a simplified setting to demonstrate how MAS works and how it compares to other methods. We additionally include a more general language modeling experiment where we explore the similarity of toxic vs nontoxic models as a case study on how to use MAS in more practical settings (Appendix B.1).
3.1 Numeric Tasks
Each task consists of a sequence of tokens that start with a beginning of sequence token, B, and end with an end of sequence token, E. Some of the tokens in each task are produced by the task environment and define the specific goals within the task. Other tokens are determined by these task provided tokens. Each trial is considered correct when all deterministic tokens are correctly predicted. During the model training, we include all token types in a Next-Token Prediction (NTP) cross entropy loss, even though many tokens are unpredictable. See Figure 3(c) for a visual depiction of each of the tasks described in this section.
@@ 3.1 Numeric Tasks
To prevent solutions that use positional readout (Grant et al., 2025), we modify the task for transformer trainings: each token in the demo phase has a 0.2 probability of being a "void" token type, V, that has no impact on the object quantity. An example sequence with an object count of 2 could be: "B V $\text{D}_{b}$ V V $\text{D}_{c}$ T R R E". All evaluations and analyses use the original Multi-Object Task.
@@ 3.2 Model Architectures
In our numeric task experiments, each model, $\psi$, is autoregressively trained to perform only one of the tasks through next-token prediction (NTP). We train 2 model seeds for each task variant up to $99.99\%$ accuracy on both training and validation data and freeze the weights before analysis and interpretation. We consider Gated Recurrent Units (GRUs) (Cho et al., 2014), Long-Short Term Memory recurrent networks (LSTMs) (Hochreiter & Schmidhuber, 1997), and two layer Transformers based on the Roformer architecture (Vaswani et al., 2017; Touvron et al., 2023; Su et al., 2023). The custom GRUs and Transformers use a dimensionality of 128, whereas the LSTM uses 64 dimensions for each the h and c vectors. We leave details of GRU and LSTM cells to the referenced papers beyond noting that the GRU and LSTM based models have the structure:
$h_{t+1}$ $=g(h_{t},x_{t})$ (4)
$\hat{x}_{t+1}$ $=f(h_{t+1})$ (5)
Where $h_{t}$ is the hidden state vector at step $t$, $x_{t}$ is the input token at step $t$, $g$ is the recurrent function (either a GRU or LSTM cell), and $f$ is a two layer (two matrix) feed-forward network (FFN) used to make a prediction, $\hat{x}_{t+1}$, of the token at step $t+1$ from the updated hidden state $h_{t+1}$.
The transformer architecture uses Rotary Positional Encodings (RoPE) (Su et al., 2023) and GELU nonlinearities (Hendrycks & Gimpel, 2023). Transformers use a history of input tokens, $X_{t}=[x_{1},x_{2},...,x_{t}]$, at each time step, $t$, to make a prediction: $\hat{h}_{t}=g(X_{t})$, and $\hat{x}_{t+1}=f(h_{t})$ where $g$ and $f$ are now sets of transformer layers. We show results from 2 layer, single attention head transformers. We refer readers to Figure 4 and Appendix B.2 for more details. For our experiments on DeepSeek-R1-Distill-Qwen-1.5B, we refer readers to Guo et al. (2025); Yang et al. (2025) for architectural details and Appendix B.1 for finetuning details.